how many kilograms of pure alcohol must be added to 20 grams of a 10% alcohol solution to produce 50%
Answer by Earlsdon(6294)

Let x = the required number of kilograms of pure alcohol.
x(100%)+(20)(10%) = (20+x)(50%) Simplify.
x+2 = 10+0.5x Subtract 0.5x from both sides.
0.5x+2 = 10 Subtract 2 from both sides.
0.5x = 8 Finally, divide both sides by 0.5
x = 16 kg.